# ScNeuGM: Scalable Neural Graph Modeling for Coloring-Based Contention and Interference Management in Wi-Fi 7

### v1

@@ V-B1 Design of Batching Mode
where $\mathcal{Q}$ represents the selected STAs and is initialized as $\emptyset$. The above process is repeated with a new query code at new random positions in each iteration until $\mathcal{Q}$ contains $K^{\prime}$ STAs. We refer to $\mathcal{Q}$ as the batch of STAs and $K^{\prime}=|\mathcal{Q}|$ as the batch size. Since the DHF is trained to output the same bits in STA hash codes if STAs are contending or hidden, as explained in Section V-A, STAs that are likely contending with and hidden from each other will likely match the query codes and be batched together. This STA batching mode is used when the NGM NN is under training to select a subset of STAs, allowing the training algorithm to focus on those contending and hidden pairs, thereby increasing efficiency.
@@ V-B2 ES with Batching Mode of DHF
To reduce the computing time of the training process of the ES, we design a curriculum learning method [46] using the batching mode of the DHF to gradually increase the number of STAs, i.e., raise the difficulty of the graph modeling task, until it reaches the total number of STAs. Specifically, as shown in Algorithm 2, the ES algorithm starts with a small number of STAs, $K^{\prime}$, and the total number of STAs, $K$. Here, the state embedding, the predictors, and the DHF are pre-trained. The distribution of the edge generator parameters is initialized with the mean $m^{\mu}_{\mathrm{EG}}=0$ and the variance $\exp({\nu^{\mu}_{\mathrm{EG}}})=\sigma^{2}_{\mathrm{0}}$. In each ES iteration, a random Wi-Fi network is simulated, and $K^{\prime}$ STAs are selected using the DHF’s batching mode. After that, the EGNN parameters are sampled from the parameterized distribution, as shown in (22), which constructs the NGM NN with the pre-trained state embedding and predictors, as shown in Fig. 5 of Section IV-A. For the given STA batch, the edges between them are generated using the NGM NN as (16), (18), and (19), forming the graph. Next, the graph is colored, and the slot assignments for the batched STAs are determined accordingly. The assignments are applied in the network, with only the batched STAs transmitting, where the transmission reliabilities are measured, and the reward is computed as in (20). Finally, the EGNN parameters’ distribution is updated in the direction indicated in (24), with $\eta$ as the learning rate. At the end of each iteration, the moving average of the previous rewards is measured as
$\Omega\leftarrow\gamma\Omega+(1-\gamma)\mathbf{1}_{\{\dot{R}\geq 0\}},$ (30)
where $\gamma$ is a smoothing factor less than 1. Since the graph’s performance is optimal when $\dot{R}\geq 0$ (or near-optimal, as an approximation of the minimum slots is used in (20)), we use $\mathbf{1}_{\{\dot{R}\geq 0\}}$ as an indicator of whether the NGM NN has satisfactory performance. In (30), if the averaged reward indicator $\Omega$ is greater than $\hat{\Omega}$, i.e., the NGM NN achieves an averaged satisfactory performance with the current number of STAs $K^{\prime}$, then the number of batched STAs $K^{\prime}$ increases by $K^{\prime\prime}$ in the next iteration; otherwise, the algorithm keeps the same number of STAs $K^{\prime}$. The algorithm terminates if it achieves an averaged satisfactory performance with $K$ STAs. The mean of the parameter distribution is returned as the trained EGNN parameters.
@@
3:  Train $\theta^{\mu}_{\mathrm{SE}}$, $\theta^{\mu}_{\mathrm{PC}}$ and $\theta^{\mu}_{\mathrm{PH}}$ as the appendix.
@@
13:     Measure reliabilities and obtain the reward $\dot{R}$ as (20).
@@ V-C STA Bucketing Mode of DHF
Furthermore, we use the DHF to bucket those highly likely contending and hidden STA pairs and generate edges only between them to reduce the computing time of the inference of the trained NGM NN.
@@ V-C1 Design of Bucketing Mode
where new random $\Psi$ positions are chosen in each repetition. Here, $\mathcal{P}$ represents the STA pairs requiring NGM NN operations and is initialized as $\emptyset$. The edges between STAs in the same bucket are generated using the NGM NN, as described in Section IV-A, while the edge values for those non-bucketed pairs are ignored and are set to 00, as
@@
Since the hash codes of contending and hidden STAs are close, they are more likely to be grouped in the same bucket, and edges between them are processed, while edges between less likely contending or hidden STAs are ignored. This STA bucketing mode is used when the trained NGM NN is deployed online across the entire network to reduce the processed edges, i.e., to reduce the complexity of the inference of the NGM NN.
V-C2 Online Graph Modeling with Bucketing Model of DHF
Figure 8: Illustration of the online graph modeling architecture.
We then present the online graph modeling architecture that deploys the NGM NN with the bucketing model of the DHF, as shown in Fig. 8. Note that we assume the NGM NN has been trained and its parameter is not updated further. In this architecture, the NGM process outputs the graph and slot assignments according to online-measured network states and runs in parallel with the network.
In the $m$-th round of the NGM process, $m=1,2,\dots$, the network measures the states $\hat{\mathcal{S}}^{(m)}$, e.g., based on the previous transmissions. Then, the NGM process uses the bucketing mode of the DHF to identify the STA pairs likely contending or interfering, indicated in $\mathcal{P}^{(m)}$, as shown in (33). Since the bucketing mode may miss some STA pairs, the process combines $\mathcal{P}^{(m)}$ with STA pairs that had edges generated in the previous $I$ rounds as
@@ V-C2 Online Graph Modeling with Bucketing Model of DHF
This is because previously generated edges are more likely to be generated in the following rounds. Here, $E_{i,j}^{(m-m^{\prime})}$ in (35) represents the edge value between STAs $i$ and $j$ in the $(m-m^{\prime})$-th round. The NGM then generates the edges between each of these pairs, forming the graph with the adjacency matrix $\mathbf{E}^{(m)}$. The slot assignments, $\mathbf{z}^{(m)}$ and $Z^{(m)}$, are determined by coloring this graph. The process takes a duration of $\tau^{(m)}$ milliseconds for the NGM process to perform the bucketing hashing, the graph modeling, and the slot assignments in total. Note that the network makes no transmissions during the first round of the NGM process, i.e., during the first $\tau^{(1)}$ milliseconds. After the slot assignments are returned from the NGM process to the network, the network starts STA transmissions with the assignments, and the NGM process begins the next round of assignments. The above process repeats until the network stops.
@@ VI-A Simulation Configurations
We use NS-3 [40] to simulate Wi-Fi networks with the RTWT mechanism. All devices are located in a 100m×$\times$100m squared area centered at the coordinates $(50,50)$ meters, simulating a factory environment [47]. We assume the network has $A=100$ APs, and they are located at the grid in the simulated area, i.e., they are at the coordinates $(5+10x,5+10y)$ meters, where $(x,y)$ are the indices of APs in $x$ and $y$ axes, i.e., $x=1,\dots,10$ and $y=1,\dots,10$. Unless specified otherwise, STAs are static, and the total number of STAs is $K=1000$. Each STA is randomly distributed within the simulated area, with each coordinate selected randomly from a uniform distribution within the interval $[0,100]$ meters. The duration of a RTWT slot is $\Delta_{\mathrm{0}}=500$ microseconds. The channel bandwidth is $B=20$ MHz at a 5.8-GHz carrier frequency. The transmission power of STAs is $\mathbf{P}_{\mathrm{0}}=0$ dBm, and the noise power is set as $\mathbb{N}_{\mathrm{0}}B=-96$ dBm. The path losses between any two devices (including all APs and STAs) follow a log-distance path loss model [48] as $-28\log_{10}(l+1)-20\log_{10}(f/(1\text{MHz}))+12$ in decibel scale, where $f$ is the carrier frequency in MHz and $l$ is the distance between two devices in meters. The receiver sensitivity threshold is set (in decibel scale) to $\mathbf{P}_{\mathrm{0}}-s_{\max}=-95$ dBm. The packet size $L$ is 800 bits. The actual decoding error probability of each transmission is computed using the same equation in (2), where $\phi$ is the signal-to-interference-plus-noise of each transmission instead. This is implemented as a transmission error model in NS-3. The target reliability of each STA is $\hat{r}=0.99$.
@@ VI-B Baseline Methods
We explain the baseline methods, other than the proposed methods. Note that since contention and interference are not measured from the network, the Markov models can hardly estimate the STA QoS. Additionally, the GNN models can hardly provide useful slot assignment decisions in graph-coloring-structured problems based on previous observations in [35] and justifications in [26, 27]. Thus, we compare the remaining approaches reviewed in Section I-A, including heuristic graph models and the NGM with PG or DPG. Their detailed implementations are explained as follows.
VI-B1 Heuristic Graph Models (IFG, CHG)
We compare the heuristic graph models used to construct binary graphs with manually designed rules. The work in [21] connects two STAs if the same AP can detect them, i.e., $E_{i,j}=\mathbf{1}_{\{\mathcal{A}_{i}\cap\mathcal{A}_{j}\neq\emptyset\}}$ for all $i\neq j$, referred to as the “IFG” scheme. Since STAs’ transmissions are primarily affected by contention and interference in the network, we also compare the graph constructed using the contention and interference indicators, where $E_{i,j}=O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j}$ for all $i\neq j$, referred to as the “CHG” scheme. Note that the CHG scheme is only applied for comparison and is impractical in real-world networks due to the unmeasured contention and interference indicators.
VI-B2 NGM Requiring Edge-Wise Feedback (PG, DPG)
We compare the designed ES algorithm with the PG [34] and DPG [36] algorithms when training the NGM NN. Specifically, the PG algorithm uses the log-derivative trick to estimate the gradient of the performance w.r.t. each edge value in (15) as
$\frac{\partial L(\theta^{\mu})}{\partial E_{i,j}}=\frac{(\dot{R}-\bar{R})\partial\log E_{i,j}}{\partial E_{i,j}},\ \forall i\neq j,$ (36)
@@ VI-C Training Convergence of SENN, PCNN, PHNN and DHF
We first show the loss values in the training process of the SENN, the PCNN and the PHNN in Fig. 9a, where the loss values are normalized against their initial values during training. Results show that the loss values decrease over the training steps and converge in less than 500 steps when training the NGM NN components, the state embedding and the predictors. Fig. 9b shows that the DHF’s losses, i.e., the similarity loss (with legend “Simi”) and the correlation loss (with legend “Corr”), converge over the iteration in Algorithm 1. Results imply that the NNs, including the state embedding, the predictors, and the DHF, are well-trained and used in the following simulations without further updates to their parameters.
@@ VI-D Configurations of Batching/Bucketing Modes of DHF
Based on the trained DHF, we then test the batching and bucketing modes. First, for the batching mode, we run the ES algorithm with a fixed small number of STAs, $K^{\prime}=20$, selected (i.e., without curriculum learning). The reward performance is measured as the moving average of the indicator $\mathbf{1}_{\dot{R}\geq 0}$ at each step. Fig. 10a compares the performance of the batching mode when using different numbers of query bits, from $\Psi=0$ to $10$, where $\Psi=0$ means the STAs are chosen randomly without using the DHF. The results indicate that the ES algorithm struggles to converge when using fewer query bits, e.g., $\Psi=0,1,2$. This is because the selected STAs are independent of their contention and interference, and they are likely to have no contention or interference between them. As a result, the NGM NN finds it difficult to learn to separate contending or hidden STAs. The convergence improves when $\Psi=3,4$. This is because more query bits lead to a higher proportion of contending or interfering STA pairs in the selected batch, which enhances training efficiency. However, the convergence slows slightly when $\Psi>4$ because the batch becomes highly biased toward contending or hidden STA pairs, and the NGM NN struggles to learn when not to separate STA pairs within the batch. Second, for the bucketing mode, we measure the proportion of hashed STA pairs from all possible STA pairs (with legend “Pair Count”), and the proportion of hashed contending or hidden pairs from all contending or hidden pairs (with legend “Recall”). Fig. 10b compares the above metrics for different numbers of query bits, from $\Psi=1$ to $15$, and hash tables, from $\Upsilon=1$ to $30$. The results show that with fewer query bits and more hash tables, more STA pairs are hashed, and the recall rate is higher. Since we aim to reduce the number of hashed STA pairs while increasing the number of hashed contending or hidden pairs among all hashed pairs, we subtract “Pair Count” from “Recall” to measure the efficiency of the bucketing mode (with legend “Efficiency”). The results indicate that $\Psi\approx 7$ and $\Upsilon\geq 20$ achieve the best efficiency. In summary, we fix $\Psi=4$ in the batching mode and $\Psi=7$ and $\Upsilon=20$ in the bucketing mode in the rest of the simulations.

### v2

@@ V-B1 Design of Batching Mode
where $\mathcal{Q}$ represents the selected STAs and is initialized as $\emptyset$. The above process is repeated with a new query code at new random positions in each iteration until $\mathcal{Q}$ contains $K^{\prime}$ STAs. We refer to $\mathcal{Q}$ as the batch of STAs and $K^{\prime}=|\mathcal{Q}|$ as the batch size. Since the DHF is trained to output the same bits in STA hash codes if STAs are contending or hidden, as explained in Section V-A, STAs that are likely contending with and hidden from each other will likely match the query codes and be batched together. This STA batching mode is used when the IGL NN is under training to select a subset of STAs, allowing the training algorithm to focus on those contending and hidden pairs, thereby increasing efficiency.
@@ V-B2 ES with Batching Mode of DHF
To reduce the computing time of the training process of the ES, we design a curriculum learning method [49] using the batching mode of the DHF to gradually increase the number of STAs, i.e., raise the difficulty of the graph-constructing task, until it reaches the total number of STAs. Specifically, as shown in Algorithm 2, the ES algorithm starts with a small number of STAs, $K^{\prime}$, and the total number of STAs, $K$. Here, the state embedding, the predictors, and the DHF are pre-trained. The distribution of the edge generator parameters is initialized with the mean $m^{\mu}_{\mathrm{EG}}=0$ and the variance $\exp({\nu^{\mu}_{\mathrm{EG}}})=\sigma^{2}_{\mathrm{0}}$. In each ES iteration, a random Wi-Fi network is simulated, and $K^{\prime}$ STAs are selected using the DHF’s batching mode. After that, the EGNN parameters are sampled from the parameterized distribution, as shown in (25), which constructs the IGL NN with the pre-trained state embedding and predictors, as shown in Fig. 5 of Section IV-A. For the given STA batch, the edges between them are generated using the IGL NN as (17), (19), and (20), forming the graph. Next, the graph is colored, and the slot assignments for the batched STAs are determined accordingly. The assignments are applied in the network, with only the batched STAs transmitting, where the transmission reliabilities are measured, and the reward is computed as in (21). Finally, the EGNN parameters’ distribution is updated in the direction indicated in (27), with $\eta$ as the learning rate. At the end of each iteration, the moving average of the previous reward performance is measured as
$\Omega\leftarrow\gamma\Omega+(1-\gamma)\mathbf{1}_{\{R\geq 0\}},$ (34)
where $\gamma$ is a smoothing factor less than 1. Since the graph’s performance is near-optimal when $R\geq 0$ (note that the maximum $R$ is greater than 0 as an approximation of the minimum slots is used in (21)(22)), we use $\mathbf{1}_{\{R\geq 0\}}$ as an indicator of whether the IGL NN has satisfactory performance. In (34), if the averaged reward indicator $\Omega$ is greater than a threshold $\hat{\Omega}$, i.e., the IGL NN achieves an averaged satisfactory performance with the current number of STAs $K^{\prime}$, then the number of batched STAs $K^{\prime}$ increases by $K^{\prime\prime}$ in the next iteration; otherwise, the algorithm keeps the same number of STAs $K^{\prime}$. The algorithm terminates if it achieves an averaged satisfactory performance with $K$ STAs. The mean of the parameter distribution is returned as the trained EGNN parameters.
@@
3: Train $\theta^{\mu}_{\mathrm{SE}}$, $\theta^{\mu}_{\mathrm{PC}}$ and $\theta^{\mu}_{\mathrm{PH}}$ as in Appendix C.
@@
13:  Measure reliabilities and obtain the reward $R$ as (21).
@@ V-C STA Bucketing Mode of DHF
Furthermore, we use the DHF to bucket those highly likely contending and hidden STA pairs and generate edges only between them to reduce the computing time of the inference of the trained IGL NN.
@@ V-C1 Design of Bucketing Mode
where new random $\Psi$ positions are chosen in each repetition. Here, $\mathcal{P}$ represents the STA pairs requiring IGL NN operations and is initialized as $\emptyset$. The edges between STAs in the same bucket are generated using the IGL NN, as described in Section IV-A, while the edge values for those non-bucketed pairs are ignored and are set to 0, as
@@
Since the hash codes of contending and hidden STAs are close, they are more likely to be grouped in the same bucket, and edges between them are processed, while edges between less likely contending or hidden STAs are ignored. This STA bucketing mode is used when the trained IGL NN is deployed online across the entire network to reduce the processed edges, i.e., to reduce the complexity of the inference of the IGL NN.
V-C2 Online Graph Construction with Bucketing Model of DHF
Figure 8: Illustration of the online graph construction architecture.
We then present the online graph construction architecture that deploys the IGL NN with the bucketing model of the DHF, as shown in Fig. 8. Note that we assume the IGL NN has been trained and its parameter is not updated further. In this architecture, the online graph construction process outputs the graph and slot assignments according to online-measured network states and runs in parallel with the network.
In the $m$-th round of the online process, $m=1,2,\dots$, the network measures the states $\hat{\mathcal{S}}^{(m)}$, e.g., based on the previous transmissions. Then, the process uses the bucketing mode of the DHF to identify the STA pairs likely contending or interfering, indicated in $\mathcal{P}^{(m)}$, as shown in (37). Since the bucketing mode may miss some STA pairs, the process combines $\mathcal{P}^{(m)}$ with STA pairs that had edges generated in the previous $I$ rounds as
@@ V-C2 Online Graph Construction with Bucketing Model of DHF
This is because previously generated edges are more likely to be generated in the following rounds. Here, $E_{i,j}^{(m-m^{\prime})}$ in (39) represents the edge value between STAs $i$ and $j$ in the $(m-m^{\prime})$-th round. The IGL then generates the edges between each of these pairs, forming the graph with the adjacency matrix $\mathbf{E}^{(m)}$. The slot assignments, $\mathbf{z}^{(m)}$ and $Z^{(m)}$, are determined by coloring this graph. The process takes a duration of $\tau^{(m)}$ milliseconds for the process to perform the bucketing hashing, the graph construction, and the slot assignments in total. Note that the network makes no transmissions during the first round of the online process, i.e., during the first $\tau^{(1)}$ milliseconds. After the slot assignments are returned from the IGL process to the network, the network starts STA transmissions with the assignments, and the process begins the next round of assignments. The above process repeats until the network stops.
@@ VI-A Simulation Configurations
We use NS-3 [43] to simulate Wi-Fi networks with the RTWT mechanism. All devices are located in a $\times100m$ squared area centered at the coordinates $(50,50)$ meters, simulating a factory environment [50]. We assume the network has $A=100$ APs, and they are located at the grid in the simulated area, i.e., they are at the coordinates $(5+10x,5+10y)$ meters, where $(x,y)$ are the indices of APs in $x$ and $y$ axes, i.e., $x=1,\dots,10$ and $y=1,\dots,10$. Unless specified otherwise, STAs are static, and the total number of STAs is $K=1000$. Each STA is randomly distributed within the simulated area, with each coordinate selected randomly from a uniform distribution within the interval $[0,100]$ meters. The duration of a RTWT slot is $\Delta_{\mathrm{0}}=500$ microseconds. The channel bandwidth is $B=20$ MHz at a 5.8-GHz carrier frequency. The transmission power of STAs is $\mathbf{P}_{\mathrm{0}}=0$ dBm, and the noise power is set as $\mathbb{N}_{\mathrm{0}}B=-96$ dBm. The path losses between any two devices (including all APs and STAs) follow a log-distance path loss model [51] as $28\log_{10}(l+1)-20\log_{10}(f/(1\text{MHz}))+12$ in decibel scale, where $f$ is the carrier frequency in MHz and $l$ is the distance between two devices in meters. The receiver sensitivity threshold is set (in decibel scale) to $\mathbf{P}_{\mathrm{0}}-s_{\max}=-95$ dBm. The packet size $L$ is 800 bits. The actual decoding error probability of each transmission is computed using the same equation in (2), where $\phi$ is the signal-to-interference-plus-noise of each transmission instead. This is implemented as a transmission error model in NS-3. The target reliability of each STA is $\hat{r}=0.99$. The maximum number of retransmissions for each STA in a RTWT slot is $5$.
@@ VI-B Baseline Methods
We explain the baseline methods, other than the proposed methods. Note that since contention and interference are not measured from the network, the Markov models can hardly estimate the STA QoS. Additionally, the GNN models can hardly provide useful slot assignment decisions in graph-coloring-structured problems based on previous observations in [37] and justifications in [25, 26]. Thus, we compare the remaining approaches reviewed in Section I-A, including human intuition-based graph models and the IGL with PG or DPG. Their detailed implementations are explained as follows.
VI-B1 Human Intuition-Based Graph Models (IFG, CHG)
We compare the human intuition-based graph models used to construct binary graphs with manually designed rules. The work in [31] connects two STAs if the same AP can detect them, i.e., $E_{i,j}=\mathbf{1}_{\{\mathcal{A}_{i}\cap\mathcal{A}_{j}\neq\emptyset\}}$ for all $i\neq j$, referred to as the “IFG” scheme. Since STAs’ transmissions are primarily affected by contention and interference in the network, we also compare the graph constructed using the contention and interference indicators, where $E_{i,j}=O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j}$ for all $i\neq j$, referred to as the “CHG” scheme. Note that the CHG scheme is applied for comparison and is impractical in the real world due to the unmeasured contention and interference indicators.
VI-B2 IGL Requiring Edge-Wise Feedback (PG, DPG)
We compare the designed ES algorithm with the PG [36] and DPG [38] algorithms when training the IGL NN. Specifically, the PG algorithm uses the log-derivative trick to estimate the gradient of the performance w.r.t. each edge value in (16) as
$\frac{\partial L(\theta^{\mu})}{\partial E_{i,j}}=\frac{(R-\bar{R})\partial\log E_{i,j}}{\partial E_{i,j}},\ \forall i\neq j,$ (40)
@@ VI-C Training Convergence of SENN, PCNN, PHNN and DHF
We first show the loss values in the training process of the SENN, the PCNN and the PHNN in Fig. 9a, where the loss values are normalized against their initial values during training. Results show that the loss values decrease over the training steps and converge in less than 500 steps when training the IGL NN components, the state embedding and the predictors. Fig. 9b shows that the DHF’s losses, i.e., the similarity loss (with legend “Simi”), the correlation loss (with legend “Corr”) and the total loss (with legend “Total”). The similarity loss converges over the iteration, while the correlation and total loss constantly decreases. We use the NNs trained at $20000$ steps, including the state embedding, the predictors, and the DHF, in the following simulations without further updates to their parameters. Note that we show the training losses in Fig. 9 to demonstrate the convergence of the training process, while the validation performance of the trained NNs is shown in the subsequent simulation results.
@@ VI-D Configurations of Batching/Bucketing Modes of DHF
Based on the trained DHF, we then test the batching and bucketing modes. First, for the batching mode, we run the ES algorithm with a fixed small number of STAs, $K^{\prime}=20$, selected (i.e., without curriculum learning). The reward performance is measured as the moving average of the indicator $\mathbf{1}_{R\geq 0}$ at each step. Fig. 10a compares the performance of the batching mode when using different numbers of query bits, from $\Psi=0$ to $10$, where $\Psi=0$ means the STAs are chosen randomly without using the DHF. The results indicate that the ES algorithm struggles to converge when using fewer query bits, e.g., $\Psi=0,1,2$. This is because the selected STAs are independent of their contention and interference, and they are likely to have no contention or interference between them. As a result, the IGL NN finds it difficult to learn to separate contending or hidden STAs. The convergence improves when $\Psi=3,4$. This is because more query bits lead to a higher proportion of contending or interfering STA pairs in the selected batch, which enhances training efficiency. However, the convergence slows slightly when $\Psi>4$ because the batch becomes highly biased toward contending or hidden STA pairs, and the IGL NN struggles to learn when not to separate STA pairs within the batch. Second, for the bucketing mode, we measure the proportion of hashed STA pairs from all possible STA pairs (with legend “Pair Count”), and the proportion of hashed contending or hidden pairs from all contending or hidden pairs (with legend “Recall”). Fig. 10b compares the above metrics for different numbers of query bits, from $\Psi=1$ to $15$, and hash tables, from $\Upsilon=1$ to $30$. The results show that with fewer query bits and more hash tables, more STA pairs are hashed, and the recall rate is higher. Since we aim to reduce the number of hashed STA pairs while increasing the number of hashed contending or hidden pairs among all hashed pairs, we subtract “Pair Count” from “Recall” to measure the efficiency of the bucketing mode (with legend “Efficiency”). The results indicate that $\Psi\approx 7$ and $\Upsilon\geq 20$ achieve the best efficiency. In summary, we fix $\Psi=4$ in the batching mode and $\Psi=7$ and $\Upsilon=20$ in the bucketing mode in the rest of the simulations.